OpenStudy (anonymous):

Help? A parallelogram whose vertices have coordinates R(1, -1), S(6, 1), T(8, 5), and U(3, 3) has a shorter diagonal of ___ . sqrt of ____

OpenStudy (anonymous):

we need a picture

OpenStudy (anonymous):

looks like short diagonal is one between the points (3,3) and (6,1) http://www.wolframalpha.com/input/?i=plot+ {%281%2C+-1%29%2C+%286%2C+1%29%2C+%288%2C+5%29%2C++%283%2C+3%29}+

OpenStudy (anonymous):

use the distance formula to find the distance \[\sqrt{(6-3)^2+(3-1)^2}=\sqrt{3^2+2^2}=\sqrt{9+4}=\sqrt{13}\]

OpenStudy (anonymous):

thank you thank you!